A math class has a total of 24 students. The number of females is 10 less than the number of males. How many males and how many females are in the class?
Answer by ikleyn(52832)

.

    m + (m-10) = 24

        2m     = 24 + 10 = 34

         m               = 34/2 = 17


ANSWER.  17 males and 17-10 = 7 females.

Solved.